OpenStudy (anonymous):

Find the value of 2x\*[\sqrt{18}x + 3x \*[\2x when x=2

OpenStudy (anonymous):

what does it mean " /* " ?

OpenStudy (anonymous):

and "\*[ "...?

OpenStudy (anonymous):

i dunno how to type it. find the value of 2x (18 under the radical sign) + 3x (2x under the radical sign)

OpenStudy (anonymous):

when x+2

OpenStudy (anonymous):

excuse me, when x=2

OpenStudy (amistre64):

you have to remove the * to get it to format; which I thought I explained :)

OpenStudy (amistre64):

\[2x\sqrt{18}x + 3x \]

OpenStudy (anonymous):

\[2x/\sqrt{18}x + 3x/2x\] ? or not?

OpenStudy (amistre64):

\*[ equation stuff goes in between here \*]

OpenStudy (amistre64):

they call it LaTeX, which is a way to tell the computer to format it in a different manner so that it looks like a math problem

OpenStudy (anonymous):

2x\*sqrt18\* + 3x \*sqrt2\* when x=2

OpenStudy (anonymous):

sorry i still havent figured it out

OpenStudy (amistre64):

it takes practice ;)

OpenStudy (amistre64):

the equation editor button on the bottom here takes most of the thinking out of it

OpenStudy (amistre64):

just type in the: \ key followed by the [ key followed by whatever you wanna type then hit the \ key followed by the ] key

OpenStudy (anonymous):

2x \[\sqrt{?}\]+ 3x\[\sqrt{?}\]when x=2

OpenStudy (amistre64):

closer ;)

OpenStudy (amistre64):

if you wanna type it all in one line: replace the [ with ( and the ] with )

OpenStudy (anonymous):

2x\[\sqrt{18}\ +3x\[\sqrt{2x}\] when x=2

OpenStudy (anonymous):

2x(\sqrt{18} + 3x\sqrt{2x when x=2

OpenStudy (amistre64):

let me see if this helps; s= ( t= ) \s 2x \sqrt{18} + 3x \sqrt{2x} \t just rep[lace s and t with their proper keystrokes

OpenStudy (anonymous):

ok sorry, Ill just have to practice with it. Thnaks a lot for your help

OpenStudy (amistre64):

:) yeah, youll get it eventually youre welcome ☺